OpenStudy (anonymous):

help me solve x2 – 4x – 8 = 0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

This is a quadratic equation that you can solve by factoring or applying quadratic formula, or completing a square method. If you can factor the equation, then that's the way to go. Otherwise, apply quadratic formula.

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

this does not factorise so u can use completing the square or the formula

OpenStudy (anonymous):

do you understand the formula?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok the general formula for a quadratic (whch u have here) is ax^2 + bx + c = 0 and u plug in the values of a, b and c into the formula which Lagrange has quoted above so in this case a = 2, b = -1 and c = -8 put these values into the formulas and u will get the solutions

OpenStudy (anonymous):

x = (-b +/- sqrt(b^2 - 4ac)) / 2a x = (-4 +/- sqrt(16 - 4*1*8)) / 2 x = (-4 +/- sqrt(-16)) / 2 x = (-4 +/- 4i) / 2 x = -2 +/- 2i

OpenStudy (anonymous):

sorry b = -4 not -1

OpenStudy (anonymous):

you've got a little wrong there lady - check your sqrt bit again

OpenStudy (anonymous):

ok thanks

OpenStudy (radar):

a=1

OpenStudy (anonymous):

oh yeah a = 1